# JavisDiT [Uncaptioned image]: Joint Audio-Video Diffusion Transformer with Hierarchical Spatio-Temporal Prior Synchronization

### v1

@@ 5.3 In-depth Analyses and Discussions
How ST-Prior Ensures Synchronized Video&Audio? Fig. 7 illustrates the synchronization mechanism for our ST-priors to guide the video-audio generation process. In particular, we visualize the spatial-temporal cross-attention map from the last block in JavisDiT at the last sampling step on both video and audio branches. The qualitative results in Fig. 7 show the spatial priors successfully help JavisDiT focus on the subject that would produce the sound (in this case, it is the bubbles rather than the diver that can make the sound), and the temporal priors bring nearly uniform attention scores along the timeline (as the bubbling sound continues from beginning to end). The cross-attention mechanism for ST-priors is well-learned for synchronized video-audio generation.
6 Related Work
In the field of AIGC, multimodal generation is an important topic, especially text-conditioned multimodal generation, such as text-to-image [41, 38, 43], text-to-video [14, 53, 61], and text-to-audio [25, 27, 16] generation, and so on. Among these tasks, the generation of sounding videos [42, 56, 51, 48, 26] has been gaining increasing attention, primarily due to its alignment with real-world application scenarios. To generate synchronized audio and video, early works decomposed the task into two cascaded subtasks. a) generating a video from text, followed by adding audio to the video in a second step [6, 19, 63, 40, 15]; or b) generating audio first, followed by synchronizing video generation with the audio [60, 18, 62]. As an instance, MovieGen [36] achieves movie-grade generation quality using this cascaded approach. From a methodological perspective, the community focus has shifted from UNet-based models [41, 27] to DiT-based methods [65, 36] to achieve more state-of-the-art performance.
Another paradigm considers the limitations of the pipeline approach, such as error propagation and reliability issues, and instead focuses on synchronized audio-video joint generation through end-to-end methods, also known as Joint Audio-Video Generation (JAVG). This paradigm emphasizes improving generation quality by modeling audio-video synchronization, including semantic consistency and temporal alignment, and has thus garnered significant attention [42, 12, 17]. For example, Ruan et al. [42] propose a diffusion block that enables interaction between the two modalities, synchronizing denoising processes. Xing et al. [56] propose using ImageBind to align the semantic representations of the two modalities. Sun et al. [48] introduce a hierarchical VAE, mapping audio and video to a shared semantic space while reducing their dimensionality. AV-DiT [51] employs a single DiT model to generate both video and audio modalities simultaneously, improving efficiency. Uniform [64] also takes a single DiT for JAVG, by simply concatenating the video and audio latent tokens during the diffusion process, without any explicit synchronization guidance. SyncFlow [26] utilizes STDiT [65] blocks to enhance the video generation quality, with a temporal adapter to guide the audio generation process. Unfortunately, current JAVG methods either lack a strong backbone for audio and video generation or insufficiently model audio-video synchronization, resulting in suboptimal generation quality. To address these issues, we propose constructing a novel DiT-based JAVG system, further enhanced by modeling hierarchical spatial-temporal synchronized prior features. Moreover, this paper aims to further advance JAVG by providing a more comprehensive, robust, and challenging benchmark.
@@ 7 Conclusion
This paper presents JavisDiT, a novel Joint Audio-Video Diffusion Transformer that simultaneously generates high-quality audio and video content with precise synchronization. We introduce the HiST-Sypo Estimator, a fine-grained spatio-temporal alignment module that extracts global and fine-grained priors to guide the synchronization between audio and video. We also propose the JavisBench dataset, comprising 10,140 high-quality text-captioned sounding videos with diverse scenes and real-world complexity, addressing limitations in current benchmarks. In addition, we introduce a temporal-aware semantic alignment mechanism to better evaluate JAVG systems on complex content. Experimental results show that JavisDiT outperforms existing approaches in both content generation and synchronization, establishing a new benchmark for JAVG tasks. Potential limitation and future work are discussed in Appendix B.
Supplementary Material
Appendix A Ethical Statement
@@
Appendix B Potential Limitation and Future Work
@@
    Scalability of Training Data: Our model was trained on 0.6M text-video-audio triplets, which, while substantial, is still limited compared to the scale of some large vision-language models (e.g., OpenSora [65] takes over 60M data to build a foundation mode for video generation). Expanding the dataset with more diverse and higher-quality real-world audio-video samples could further enhance the model’s ability to generalize across different domains and fine-grained synchronization patterns.
@@
    Efficiency and Computation Overhead: JavisDiT employs a Diffusion Transformer for high-quality generation, but diffusion-based models tend to be computationally intensive. While our approach achieves state-of-the-art results, generation speed and efficiency remain challenges (e.g., generating a 2-second sounding-video at 720P/24fps/16kHz takes 6 minutes on one H100 GPU). Exploring accelerated sampling strategies or hardware optimization could improve efficiency.
@@
    Benchmarking Across Resolutions and Durations: Our evaluations primarily focus on a fixed resolution (240P) and duration (4s) setting in JavisBench. However, real-world applications may require generation at higher resolutions (e.g., 720P, 1080P) and longer durations. Conducting benchmark tests across multiple settings would provide a more comprehensive understanding of current models’ strengths and limitations, benefiting the research community with standardized evaluations.
@@
By addressing these limitations, future iterations of JavisDiT could further enhance scalability, synchronization accuracy, efficiency, and adaptability, pushing the boundaries of joint audio-video generation.
Appendix C Implementation Details
C.1 JavisDiT Model Configuration
Model Architecture. As Fig. 2 illustrates, JavisDiT consists of two branches for video and audio generation, each comprising $N=28$ DiT blocks. Within each DiT block, the latent video and audio representations are processed through several modules, where all attention modules utilize 16 attention heads with a hidden size of 1152. The intermediate dimension of the FFN is 4×4$\times$4 × the hidden size. Each attention and FFN module is preceded by a LayerNorm layer. In practical implementation, every attention module is followed by an FFN module to further enhance feature modeling. The video/audio latent is sequentially processed with: Spatial-SelfAttnCrossAttnSpatial-CrossAttnBi-CorssAttnFFNTemporal-SelfAttnCrossAttnTemporal-CrossAttnBi-CorssAttnFFN for $N=28$ times, resulting in our JavisDiT with 3.14B parameters in total.
@@ C.1 JavisDiT Model Configuration
Figure A1: Stage1: audio pretraining. Parameters are initialized from the video branch.
Training Strategy. As Sec. 3.3 states, we carefully design a three-stage training strategy to achieve high-quality single-modal generation and ensure fine-grained spatio-temporal synchrony on generated videos and audios. In particular, the video branch of JavisDiT (including the ST-SelfAttn module, the Coarse-Graind CrossAttn module, and the FFN module) is initialized from OpenSora [65] and frozen during the whole training stages. We use the weights of the video branch to initialize the audio branch for better convergence (see Fig. A1). The detailed configuration for the three-stage training strategy is displayed in Tab. A1.
Training Data Curation. For audio-pretraining (stage-1), the audio-caption pairs come from various public datasets, including AudioSet [10], AudioCaps [20], VGGSound [2], WavCaps [33], Clotho [7], ESC50 [35], GTZAN [47], MACS [32], and UrbanSound8K [44]. They contribute to a total of 788K training entries. For the ST-Prior estimator and the final JavisDiT training (stage-2&3), the data source comes from two newly-proposed sounding video datasets: MMTrail [4] and TAVGBench [31]. MMTrail provides slightly more caption annotations (2M) than TAVGBench (1.7M) but has lower quality. Due to the resource limit, we collect a part of the original sounding-videos from YouTube and filtered out around 80% human-talking videos with the FunASR  [^2] https://github.com/modelscope/FunASR tool. We finally collect 136K high-resolution video-audio-caption triplets from MMTrail and 475K from TAVGBench, resulting in 611K data for training in total. The data construction for ST-Prior estimator’s contrastive learning is discussed in Sec. C.2.4.
Training Objective and Inference. The ST-Prior Estimator is trained with contrastive learning objectives, which will be detailed in Sec. C.2. For the DiT model, we use rectified flow [29] as the denoising scheduler for better performance [65, 36]. The inference sample step is 30, with classifier-free guidance at 7.0. The video and audio latent are concurrently sampled at each inference step. The video encoder-decoder comes from OpenSora [65], and the audio encoder-decoder comes from AudioLDM2 [27]. Both of them are frozen during training.
Detailed Evaluation Setup. In Sec. 5 in our manuscript, we conduct a comprehensive evaluation across multiple video and/or audio generation models on two settings: (1) our proposed JavisBench benchmark and (2) previously used AIST++ [22] and Landscape [21] datasets. For our JavisBench, generative models are required to generate videos and 240p and 24fps, with audios at 16kHz sample rate. The duration is 4 seconds. For AIST++ [22], we follow the literature [42, 48] to generate 2-second clips at the visual resolution of $256\times 256$, with 240p on landscape [21].
@@
epoch 13 1 2
training objective rectified flow contrastive rectified flow
GPU days (H100) 16 8 256
@@ C.2.1 Overview
Since ImageBind [11] has built a consistent semantic space across multiple modalities, we use ImageBind’s text encoder to extract the potential spatio-temporal prior from the input text caption. Formally, for a given input text $s$, we leverage the 77 hidden states from the last layer of ImageBind’s text encoder, and learn $N_{s}$ spatial tokens $\mathbf{p}_{s}$ and $N_{t}$ temporal tokens $\mathbf{p}_{t}$ from a 4-layer transformer block $\mathcal{P}$. Notably, since the same event may occur at different locations and timestamps in different video-audio generations, the same text $s$ can also produce different ST-Priors $(\mathbf{p}_{s},\mathbf{p}_{t})$. We adopt a sampling strategy akin to VAE, where our ST-Prior Estimator $\mathcal{P}$ outputs the mean and variance of a Gaussian distribution, from which we sample a plausible $(\mathbf{p}_{s},\mathbf{p}_{t})$ to describe specific spatio-temporal events. This can be formalized as: $(\mathbf{p}_{s},\mathbf{p}_{t})=\mathcal{P}_{\phi}(s;\epsilon)$, where $\epsilon$ refers to the normalized Gaussian distribution.
@@
Formally, we also extract $N_{s}$/$N_{t}$ positive or negative spatial/temporal embeddings from synchronous or asynchronous video-audio pairs: $(\mathbf{e}_{s}^{+},\mathbf{e}_{t}^{+})$=$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a})$, $(\mathbf{e}_{s}^{-},\mathbf{e}_{t}^{-})$=$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a}^{-})\;\text{or}\;\mathcal{E}_{\psi}(\mathbf{v}^{-},\mathbf{a})$. Sec. C.2.2 displays more details.
The training objective is to make the priors $(\mathbf{p}_{s},\mathbf{p}_{t})$ closer to the synchronous embeddings $(\mathbf{e}_{s}^{+},\mathbf{e}_{t}^{+})$ while pushing them further from the asynchronous $(\mathbf{e}_{s}^{-},\mathbf{e}_{t}^{-})$, thereby equipping the priors $(\mathbf{p}_{s},\mathbf{p}_{t})$ with fine-grained conditioning capabilities. Several contrastive losses with the Kullback-Leibler (KL) divergence loss are utilized to optimize the spatial and temporal priors (more details refer to Sec. C.2.3):
$\mathcal{L}_{prior}=$ $\mathcal{L}_{\text{contrast}}(\mathbf{p}_{s},\mathbf{e}_{s}^{+},\mathbf{e}_{s}^{-})+\mathcal{L}_{\text{kl}}(\mathbf{p}_{s})+$ (A1)
$\mathcal{L}_{\text{contrast}}(\mathbf{p}_{t},\mathbf{e}_{t}^{+},\mathbf{e}_{t}^{-})+\mathcal{L}_{\text{kl}}(\mathbf{p}_{t})\,.$

### v2

@@ 5.3 In-depth Analyses and Discussions
How does ST-Prior ensure synchronized video&audio? Fig. 7 illustrates the synchronization mechanism for our ST-priors to guide the video-audio generation process. In particular, we visualize the spatial-temporal cross-attention map from the last block in JavisDiT at the last sampling step on both video and audio branches. The qualitative results in Fig. 7 show the spatial priors successfully help JavisDiT focus on the subject that would produce the sound (in this case, it is the bubbles rather than the diver that can make the sound), and the temporal priors bring nearly uniform attention scores along the timeline (as the bubbling sound continues from beginning to end). Sec. E.5 also details how spatiotemporal priors are injected and shifted over sequential events, further illustrating the well-learned cross-attention mechanism of ST-priors for synchronized video-audio generation.
JavisDiT supports variant-length sounding video generation. In the previous sections, our evaluations primarily focused on generating 4-second videos. Here, we extend the assessment to 10-second outputs and report several key metrics. The experimental results in Tab. 5 show that our JavisDiT maintains strong visual and auditory quality (FVD, FAD), semantic consistency (CLIP, CLAP, AVHScore), and audio–video synchrony (JavisScore) when generating 10-second videos. These findings provide strong evidence that our approach effectively supports diverse audio–video generation requirements across different scenarios, durations, and resolutions.
Table 5: Quantitative evaluation on variant-length generation performance.
[…]
4s 241.8 7.3 0.308 0.382 0.186 0.153
10s 233.8 7.1 0.307 0.385 0.183 0.154
5.4 Human Evaluation
In this section, we further conduct user studies to more comprehensively evaluate the performance of joint audio–video generation through human evaluation. Specifically, we randomly sample 100 prompts from JavisBench and run UniVerse-1 (Wang et al., 2025) and our JavisDiT to generate corresponding audio–video outputs. Three volunteers are then recruited to perform blind win–tie–lose preference judgments on three aspects: video quality, audio quality, and audio-video alignment. The averaged scores across annotators are used as the final evaluation metric.
Figure 8: Human evaluation on generation performance.
The human evaluation results in Fig. 8 are largely consistent with the objective metrics reported in Tab. 1:
(1) Our video quality is indeed slightly lower than that of UniVerse-1 (Wang et al., 2025). As analyzed in Sec. 5.2, this mainly stems from the performance gap between the backbones: our method currently relies on the OpenSora-1.2 (Zheng et al., 2024) backbone, which is notably weaker than the Wan-2.1 (Wan et al., 2025) backbone used in UniVerse-1. We expect this issue to be substantially alleviated as stronger T2V backbones become available in future versions.
(2) Our audio quality and audio–video alignment outperform UniVerse-1. This is because UniVerse-1 adopts a relatively complex architecture to bridge two separately pretrained branches, and the resulting representation gap weakens cross-modal interaction. In contrast, our JavisDiT design offers a more streamlined and effective mechanism for enabling tight audio–video coupling.
@@ 6 Conclusion
This paper presents JavisDiT, a novel Joint Audio-Video Diffusion Transformer that simultaneously generates high-quality audio and video content with precise synchronization. We introduce the HiST-Sypo Estimator, a fine-grained spatio-temporal alignment module that extracts global and fine-grained priors to guide the synchronization between audio and video. We also propose the JavisBench dataset, comprising 10,140 high-quality text-captioned sounding videos with diverse scenes and real-world complexity, addressing limitations in current benchmarks. In addition, we introduce a temporal-aware semantic alignment mechanism to better evaluate JAVG systems on complex content. Experimental results show that JavisDiT outperforms existing approaches in both content generation and synchronization, establishing a new benchmark for JAVG tasks. Potential limitations and future work are discussed in Sec. A.5.
Appendix A Discussions
A.1 Ethical Statement
@@
A.2 Societal Impact Statement
This work proposes JavisDiT, a unified model for synchronized audio-video generation from open-ended prompts. While primarily intended as a contribution to foundational multimodal generation research, it has both positive applications and potential societal risks.
Positive Impact. JavisDiT may support creative industries by enabling efficient generation of audio-video content. Addressing the challenge of audio-video alignment has practical implications for animation, video conferencing, television broadcasting, and video editing—domains where synchronization is traditionally achieved through offline computation or extensive manual post-processing.
Potential Risks. As with other generative models, JavisDiT may be misused to produce highly realistic synthetic media for malicious purposes, including misinformation, impersonation, and manipulation. Its ability to generate temporally coherent sounding videos may increase the realism and persuasiveness of deepfakes. Specifically, fabricated audio-video clips could be used to spread false information or simulate events that never occurred, posing risks to public trust, political discourse, and media integrity.
Mitigation Strategy. To mitigate these concerns, we recommend responsible release practices, including usage gating and dataset transparency. Future work may explore watermarking and detection tools to distinguish synthetic output.
A.3 Reproducibility Statement
We provide detailed descriptions of model design, training, and evaluation in both the main paper and the appendix. Furthermore, all code, pretrained checkpoints, and processed datasets will be publicly released to ensure full reproducibility of our results.
A.4 LLM Usage Statement
Large Language Models (LLMs) were used solely as writing assistants, including tasks such as language polishing and presentation refinement. They were not involved in the conception of core ideas or designs.
A.5 Potential Limitation and Future Work
@@
    Scalability of Training Data: Our model was trained on 0.6M text-video-audio triplets, which, while substantial, is still limited compared to the scale of some large vision-language models (e.g., OpenSora (Zheng et al., 2024) takes over 60M data to build a foundation mode for video generation). Expanding the dataset with more diverse and higher-quality real-world audio-video samples could further enhance the model’s ability to generalize across different domains and fine-grained synchronization patterns.
@@
    Efficiency and Computation Overhead: JavisDiT employs a Diffusion Transformer for high-quality generation, but diffusion-based models tend to be computationally intensive, as analyzed in Tab. A1. While our approach achieves state-of-the-art results, generation speed and efficiency remain challenges (e.g., generating a 2-second sounding-video at 720P/24fps/16kHz takes 6 minutes on one H100 GPU). Exploring accelerated sampling strategies or hardware optimization could improve efficiency.
@@
    Benchmarking Across Resolutions and Durations: Our evaluations primarily focus on a fixed resolution (240P) and duration (4s) setting in JavisBench. However, real-world applications may require generation at higher resolutions (e.g., 1080P) and longer durations. Conducting benchmark tests across multiple settings would provide a more comprehensive understanding of current models’ strengths and limitations.
Table A1: Latency analysis for 240P, 4s sounding-video generation on H100 with 30 sampling steps.
Video Branch Audio Branch ST-Prior Modulation AV-Interaction Overall
13s 2s 13s 2s 30s
@@
By addressing these limitations, future iterations of JavisDiT could further enhance scalability, synchrony, efficiency, and adaptability, pushing the boundaries of joint audio-video generation.
Appendix B Related Work
In the field of AIGC, multimodal generation has become a key topic, encompassing text-to-image (Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022), text-to-video (Ho et al., 2022; Wang et al., 2024d; Yu et al., 2023), and text-to-audio (Liu et al., 2023a; 2024b; Huang et al., 2023) generation, etc.. Among these, sounding video generation (Ruan et al., 2023; Xing et al., 2024; Wang et al., 2024b; Sun et al., 2024; Liu et al., 2024a) is drawing increasing attention due to its strong alignment with real-world applications. To generate synchronized audio and video, early works decomposed the task into two cascaded subtasks. a) generating video from text, then adding audio (Comunità et al., 2024; Jeong et al., 2024; Zhang et al., 2024; Ren et al., 2024; Hu et al., 2024); or b) generating audio first, then synchronizing video (Yariv et al., 2024; Jeong et al., 2023; Zhang et al., 2025). As an instance, MovieGen (Polyak et al., 2024) achieves movie-grade generation quality using this cascaded approach. From a methodological perspective, the community focus has shifted from UNet-based models (Rombach et al., 2022; Liu et al., 2024b) to DiT-based methods (Zheng et al., 2024; Polyak et al., 2024) to achieve more state-of-the-art performance.
Another line of work considers the limitations of pipeline approaches, such as error propagation and reliability issues, by focusing on end-to-end Joint Audio-Video Generation (JAVG). This paradigm aims to improve generation quality by modeling audio-video synchronization, including semantic consistency and temporal alignment, and has received growing attention (Ruan et al., 2023; Hayakawa et al., 2024; Ishii et al., 2024). For example, Ruan et al. (2023) introduce a diffusion block that enables cross-modal interaction during denoising. Xing et al. (2024) propose using ImageBind to align the semantic representations of the two modalities. Sun et al. (2024) introduce a hierarchical VAE, mapping audio and video to a shared semantic space while reducing their dimensionality. AV-DiT (Wang et al., 2024b) employs a single DiT model to generate both video and audio modalities simultaneously, improving efficiency. Uniform (Zhao et al., 2025) also takes a single DiT for JAVG, by simply concatenating the video and audio latent tokens during the diffusion process, without any explicit synchronization guidance. SyncFlow (Liu et al., 2024a) utilizes STDiT (Zheng et al., 2024) blocks to enhance the video generation quality, with a temporal adapter to guide the audio generation process. Unfortunately, current JAVG methods either lack a strong backbone for audio and video generation or insufficiently model audio-video synchronization, resulting in suboptimal generation quality. To address these issues, we construct a novel DiT-based JAVG model, further enhanced by modeling hierarchical spatial-temporal synchronized prior features. Moreover, this paper aims to further advance JAVG by providing a more comprehensive, robust, and challenging benchmark.
@@ Appendix B Related Work
Figure A1: Audio Pretraining. Parameters are initialized from the video branch.
@@
epoch 55 1 2
training objective rect. flow contrastive rect. flow
GPU days (H100) 64 8 256
Appendix C Implementation Details
C.1 JavisDiT Model Configuration
Model Architecture. As Fig. 2 illustrates, JavisDiT consists of two branches for video and audio generation, each comprising $N=28$ DiT blocks. Within each DiT block, the latent video and audio representations are processed through several modules, where all attention modules utilize 16 attention heads with a hidden size of 1152. The intermediate dimension of the FFN is $4\times$ the hidden size. Each attention and FFN module is preceded by a LayerNorm layer. In practical implementation, every attention module is followed by an FFN module to further enhance feature modeling. The video/audio latent is sequentially processed with: Spatial-SelfAttnCrossAttnSpatial-CrossAttnBi-CorssAttnFFNTemporal-SelfAttnCrossAttnTemporal-CrossAttnBi-CorssAttnFFN for $N=28$ times, resulting in our JavisDiT with 3.14B parameters in total.
In particular, we adopt spatial–temporal (self- or cross-) attention modules primarily to reduce computational cost. For video tokens of size ($T\times H\times W$), the computational complexity of full self-attention is $O\big((THW)^{2}\big)$. In contrast, spatial–temporal attention reduces the complexity to $O\big(T(HW)^{2}+T^{2}(HW)\big)=O\big(THW\cdot(T+HW)\big)$, which substantially lowers both memory consumption and computational overhead.
Training Strategy. As Sec. 3.3 states, we carefully design a three-stage training strategy to achieve high-quality single-modal generation and ensure fine-grained spatio-temporal synchrony on generated videos and audios. In particular, the video branch of JavisDiT (including the ST-SelfAttn module, the Coarse-Graind CrossAttn module, and the FFN module) is initialized from OpenSora (Zheng et al., 2024) and frozen during the whole training stages. We use the weights of the video branch to initialize the audio branch for better convergence (see Tab. A2). The detailed configuration for the three-stage training strategy is displayed in Tab. A2.
Training Data Curation. For audio-pretraining (stage-1), the audio-caption pairs come from various public datasets, including AudioSet (Gemmeke et al., 2017b), AudioCaps (Kim et al., 2019), VGGSound (Chen et al., 2020a), WavCaps (Mei et al., 2024), Clotho (Drossos et al., 2020), ESC50 (Piczak, 2015), GTZAN (Sturm, 2013), MACS (Martín-Morató & Mesaros, 2021), and UrbanSound8K (Salamon et al., 2014). They contribute to a total of 788K training entries. For the ST-Prior estimator and the final JavisDiT training (stage-2&3), the data source comes from two newly-proposed sounding video datasets: MMTrail (Chi et al., 2024) and TAVGBench (Mao et al., 2024). MMTrail provides slightly more caption annotations (2M) than TAVGBench (1.7M) but has lower quality. Due to the resource limit, we collect a part of the original sounding-videos from YouTube and filtered out around 80% human-talking videos with the FunASR  [^2] https://github.com/modelscope/FunASR tool. We finally collect 136K high-resolution video-audio-caption triplets from MMTrail and 475K from TAVGBench, resulting in 611K data for training in total. Fig. A2 visualizes the prompt distribution, which further demonstrates the diversity of our collected training dataset. The data construction for ST-Prior estimator’s contrastive learning is discussed in Sec. C.2.4.
Figure A2: T-SNE visualization of prompt distribution of the training dataset.
Training Objective and Inference. The ST-Prior Estimator is trained with contrastive learning objectives, which will be detailed in Sec. C.2. For the DiT model, we use rectified flow (Liu et al., 2023b) as the denoising scheduler for better performance (Zheng et al., 2024; Polyak et al., 2024). The inference sample step is 30, with classifier-free guidance at 7.0. The video and audio latent are concurrently sampled at each inference step. The video encoder-decoder comes from OpenSora (Zheng et al., 2024), and the audio encoder-decoder comes from AudioLDM2 (Liu et al., 2024b). Both of them are frozen during training.
Detailed Evaluation Setup. In Sec. 5 in our manuscript, we conduct a comprehensive evaluation across multiple video and/or audio generation models on two settings: (1) our proposed JavisBench benchmark and (2) previously used AIST++ (Li et al., 2021) and Landscape (Lee et al., 2022) datasets. For our JavisBench, generative models are required to generate videos and 240p and 24fps, with audios at 16kHz sample rate. The duration is 4 seconds. For AIST++ (Li et al., 2021), we follow the literature (Ruan et al., 2023; Sun et al., 2024) to generate 2-second clips at the visual resolution of $256\times 256$, with 240p on landscape (Lee et al., 2022).
@@ C.2.1 Overview
Since ImageBind (Girdhar et al., 2023) has built a consistent semantic space across multiple modalities, we use ImageBind’s text encoder to extract the potential spatio-temporal prior from the input text caption. Formally, for a given input text $s$, we leverage the 77 hidden states from the last layer of ImageBind’s text encoder, and learn $N_{s}$ spatial tokens $\mathbf{p}_{s}$ and $N_{t}$ temporal tokens $\mathbf{p}_{t}$ from a 4-layer transformer block $\mathcal{P}$. Notably, since the same event may occur at different locations and timestamps in different video-audio generations, the same text $s$ can also produce different ST-Priors $(\mathbf{p}_{s},\mathbf{p}_{t})$. We adopt a sampling strategy akin to VAE, where our ST-Prior Estimator $\mathcal{P}$ outputs the mean and variance of a Gaussian distribution, from which we sample a plausible $(\mathbf{p}_{s},\mathbf{p}_{t})$ to describe specific spatio-temporal events. This can be formalized as: $(\mathbf{p}_{s},\mathbf{p}_{t})=\mathcal{P}_{\phi}(s;\epsilon)$, where $\epsilon$ refers to the normalized Gaussian distribution.
@@
Formally, we also extract $N_{s}$$N_{t}$ positive or negative spatial/temporal embeddings from synchronous or asynchronous video-audio pairs: $(\mathbf{e}_{s}^{+},\mathbf{e}_{t}^{+})$$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a})$, $(\mathbf{e}_{s}^{-},\mathbf{e}_{t}^{-})$$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a}^{-})\;\text{or}\;\mathcal{E}_{\psi}(\mathbf{v}^{-},\mathbf{a})$. Sec. C.2.2 displays more details.
The training objective is to make the priors $(\mathbf{p}_{s},\mathbf{p}_{t})$ closer to the synchronous embeddings $(\mathbf{e}_{s}^{+},\mathbf{e}_{t}^{+})$ while pushing them further from the asynchronous $(\mathbf{e}_{s}^{-},\mathbf{e}_{t}^{-})$, thereby equipping the priors $(\mathbf{p}_{s},\mathbf{p}_{t})$ with fine-grained conditioning capabilities. Several contrastive losses with the Kullback-Leibler (KL) divergence loss are utilized to optimize the spatial and temporal priors (see Sec. C.2.3):
$\mathcal{L}_{prior}=$ $\mathcal{L}_{\text{contrast}}(\mathbf{p}_{s},\mathbf{e}_{s}^{+},\mathbf{e}_{s}^{-})+\mathcal{L}_{\text{kl}}(\mathbf{p}_{s})+\mathcal{L}_{\text{contrast}}(\mathbf{p}_{t},\mathbf{e}_{t}^{+},\mathbf{e}_{t}^{-})+\mathcal{L}_{\text{kl}}(\mathbf{p}_{t})\,.$ (A1)